How do I solve this equation, please show steps.
4x - 5(3x + 10) = 126
I've tried distrubuting 5 to 3x and 10 to get 15x + 50. So then my equation is 4x - 15 + 50. So I add 15 to itself crossing it out, then I add it to 50 which is 65. Im stuck on this part I dont know what to do.

Found 3 solutions by user_dude2008, stanbon, rfer:
Answer by user_dude2008(1862)

4x - 5(3x + 10) = 126
4x - 15x - 50 = 126
4x - 15x = 126 +50
-11x = 176
x = 176/(-11)
x=-16 <---------------- Answer

Answer by stanbon(75887)

4x - 5(3x + 10) = 126
4x - 15x - 50 = 126
Simplify the left side:
-11x - 50 = 126
Add 50 to both sides:
-11x = 176
Divide both sides by -11 to get:
x = -16
=============
Cheers,
Stan H.

Answer by rfer(16322)

4x-5(3x+10)=126
4x-15x-50=126
add 50 and combine like quanities
-11x=176
divide by -11
x=-16